OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -2 f(1) = 8 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? (4 points) Question 4 options: 1) -2, 8, -26, -80, … 2) -2, 8, -26, 80, … 3) -2, 12, -44, 180, … 4) -2, -12, -44, -180, …

OpenStudy (anonymous):

Its either option A or B

OpenStudy (anonymous):

f(2)=-4 f(1)-3 f(0) = -32+6 = -26 f(3)=-4 f(2) - 3f(1) = 104 - 24 = 80

OpenStudy (anonymous):

so its B